OpenStudy (anonymous):

find the equation of the line tangent to the graph y=sqrt(x) at x=9 help pls

OpenStudy (anonymous):

\[\Large \sqrt{x}=x^{\frac{1}{2}}\]

OpenStudy (anonymous):

nevermind i think i got it

OpenStudy (anonymous):

The power rule basically states \[\Large \frac{ d }{ dx }(x^{n})=nx^{n-1}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

would it be dy/dy=1/2(x)?

OpenStudy (anonymous):

i meant dy/dx=1/2sqrt(x)